OpenStudy (anonymous):

During an intramural basketball game, Team A scored a total of 142 points. How many points did Team A score during the game.

OpenStudy (anonymous):

the question you ask states the answer. 142 points. with the way you worded the question that is the answer. Is there anymore information?

OpenStudy (anonymous):

my teacher said 63 but I cant get how to solve it @smashinsam00

OpenStudy (ash2326):

@nia_maepenix Would you please check the question again? in case some info is missing:)

OpenStudy (anonymous):

During an intramural basketball game, Team A 16 fewer points than Team B. Together, both teams scored a total of 142 points. How many points did Team A score during the game. @ash2326

OpenStudy (ash2326):

Let team B score x points then team A would have scored x-16 points

OpenStudy (ash2326):

do you get this?

OpenStudy (anonymous):

x+x-16=142? @ash2326

OpenStudy (ash2326):

yes, now could you find x from this?

OpenStudy (anonymous):

you can set up the equation like: x+ (x-16) = 142 and solve for x

OpenStudy (anonymous):

2x=142+16=158

OpenStudy (anonymous):

since the number of points team A scored is (x-16) fill in x with what you found it to be

OpenStudy (anonymous):

x=79 and x-16=63

OpenStudy (ash2326):

great work @nia_maepenix :)

OpenStudy (anonymous):

Thanks guys!